OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (anonymous):

@asnaseer @PFEH.1999

OpenStudy (anonymous):

to find the inverse: replace f(x) by y: y = x^2 - 16 now invert x and y x = y^2 - 16. now put it back in the form of f(x) = y y = sqrt(x + 16)

OpenStudy (anonymous):

the new y is f-1(x)

OpenStudy (anonymous):

\[f⁻¹(y) = \sqrt{y + 16}\]

OpenStudy (anonymous):

here's the answer choices though so would it be B?

OpenStudy (anonymous):

yes, B

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

glad to help :)